a train cover certain distance between two places at uniform speed if train move 10km/hr faster it will take 2 hours less.if train move 10km/hr slower it will take 3 hours more than the scheduled time what is distance covered by train
Answer by mananth(16946)

a train cover certain distance between two places at uniform speed if train move 10km/hr faster it will take 2 hours less.if train move 10km/hr slower it will take 3 hours more than the scheduled time what is distance covered by train
let distance =x
and rate for scheduled time be r
let time be t
x=rt
x/(r+10)=t-2
x=(t-2)(r+10)
x/(r-10)=t+3
x=(t+3)(r-10)
(t-2)(r+10)=rt
tr+10t-2r-20=rt
10t-2r=20....................(1)
(t+3)(r-10)=rt
tr-10t+3r-30=rt
-10t+3r=30.....................(2)
add (1) & (2)
r=50
plug r in (1)
10t-100=20
10t=120
/10
t=12
x=rt
x=600 miles